OpenStudy (anonymous):

cos 24degree + cos 5 degree+ cos 175degre+ cos 204degree+cos 300 degree =

OpenStudy (anonymous):

Just rewrite them, 175 = 180-5 , 204 = 180 +24, 300 = 365-300 etc..